c. y + 3y + 2y = sin(e) hint: sub yx d. x²y – 2xy + 4y = 0 into robe
Answer #1

Given yl+3y+ 24= Sinter. Auxilliary-eavaton associatea to Romogeno us defp. easchm Solution. > m2+2m +m+2=0 m2+3m+2=0 Cm+2)(yox)= 2X Siner + Gēx+zē? Aus. 22y11-22ey+4y=0 Put x= e2 2= logae. yl= d4.dz =) Deyl=dy dz. Similasty 2e2yllc cRy d22 Then eau